The level of thorium in a sample decreases by a factor of one-half every 4.2 million years. A meteorite is discovered to have only 7.6% of it's original thorium remaining. How old is the meteorite?: The level of thorium in a sample decreases by a factor of one-half every 4.2 million years. A meteorite is discovered to have only 7.6% of it's original thorium remaining. How old is the meteorite?
Answer by ankor@dixie-net.com(4489)

The level of thorium in a sample decreases by a factor of one-half every 4.2 million years. A meteorite is discovered to have only 7.6% of it's original thorium remaining. How old is the meteorite?
:
The half-life formula:
A = Ao*2^(-t/h)
where:
Ao = initial amt (= 1 in this problem)
A = resulting amt (decimal equiv of %)
t = time (in millions of years here)
h = half-life of the substance (millions of years)
:
.076 = 1*2^(-t/4.2)
or
2^(-t/4.2) = .076
:
Find the nat log of both sides:
(-t/4.2)*.693147 = -2.577022
:
-t/4.2 = -2.577022/.693147
Cross multiply:
-.693147t = 4.2 * -2.577022
:
-.693147t = -10.8235
t = 10.8235/.693147; (minus into a minus is plus)
t = 15.615 million years old
:
:
Check solution on a calc: enter 2^(-15.615/4.2) = .075999 ~ .076 or 7.6%